OpenStudy (anonymous):

How do I show my work for this? I have the answer but I need to show my work. 4x/9 + x-10/9 = 5(x-2)/9

OpenStudy (anonymous):

\[\frac{ 4x }{ 9 }+x-\frac{ 10 }{9 }=\frac{ 5\left( x-2 \right) }{ 9 }\] multiply by 9, 4x+9x-10=5x-10 4x+9x-5x=-10+10 8x=0 x=0

OpenStudy (anonymous):

(t²+6t)/(t-1) - 7/(t-1) = (x+7)/1 @surjithayer can you help with this one also I have another four if you can. Please and thank you

OpenStudy (anonymous):

\[\frac{ t ^{2}+6t }{ t-1}-\frac{ 7 }{ t-1 }=\frac{ x+7 }{ 1 }\] \[\frac{t ^{2}+6t-7 }{t-1 }=\frac{ x+7 }{1 }\] \[\frac{ t ^{2}+\left( 7-1 \right)t-7 }{t-1 }=\frac{ x+7 }{ 1 }\] \[\frac{ t ^{2}+7t-t-7 }{t-1}=\frac{ x+7 }{1 }\] \[\frac{ t \left( t+7 \right)-1\left( t+7 \right) }{ t-1 }=\frac{ x+7 }{ 1 }\] \[\frac{ \left( t+7 \right)\left( t-1 \right) }{t-1 }=x+7\] t+7=x+7 t=x